OpenStudy (anonymous):

2x^4 + 5x^3 + 3x^2 + 15x – 9 = 0 Find all roots and factor completely

OpenStudy (solomonzelman):

2x^4 + 5x^3 + 3x^2 + 15x – 9 = 0 (x+3)(2x-1)(x^2+3)=0 x=1/2 or x=-3

OpenStudy (solomonzelman):

Did this help you?

OpenStudy (solomonzelman):

can you please give me a medal (by clicking a best response)? Please!